If money supply increased by 2%, real output increased by 1%, and prices increased by 1%, is money neutral? Why?

2. According to the quantity of money, if money stock is equal to 10 trillion euro, real GDP is 30 trillion euro, and price index is 1.2, what is V?

3. How do you interpret the number you calculated in the previous subquestion?

4. In your job interview for a consulting position with Chase Bank, you receive the following data for one of the countries in which Chase operates. In 2016, money supply was 20 trillion, in 2017, it grew to 21 trillion. Price level grew 8% over the same time period. The interviewer tells you that it is safe to use the regular assumption about velocity of money in that economy and asks you to calculate the real output growth rate and the nominal output growth rate between 2017 and 2016.

5. For a hefty starting salary bonus, the Chase interviewer asks you to derive the equation that you used to give them the answer to the previous subquestion. However, do not omit any steps.
Answer #1

1)

Money is neutral where rise in money supply does not affect real variables. Real output is real variable. Hence, money supply is affecting real output over here, thus, Money is not neutral here.

2 % rise in money supply lead to rise in the output by 1 %.

2)

MV =PY

10*V = 1.2*30

10*V = 36

V = 36/10

= 3.6